OpenStudy (anonymous):

divide, simplify 4p-4 over p divided by 7p-7 over 8p^2

OpenStudy (anonymous):

\[\frac{\frac{4p-4}{p}}{\frac{7p-7}{8p^2}} = \frac{4p-4}{p} \times \frac{8p^2}{7p-7}\]

OpenStudy (anonymous):

\[\frac{(4p-4)(8p^2)}{(p)(7p-7)} = \frac{32p^3-32p^2}{7p^2-7p} = \frac{32p(p-1)}{7(p-1)} = \frac{32p}{7}\]

OpenStudy (anonymous):

you're a lifesaver, ty

OpenStudy (anonymous):

yw